OpenStudy (anonymous):

for y=2/3x-2 where the slope is 2/3 and the y- intercept is -2 can someone help me Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences. ?

OpenStudy (anonymous):

They already give you the line in slope-intercept method. For graphing it you know the y intercept is -2 so put y=-2 on x=0 and just use some other coordinate

OpenStudy (anonymous):

wouldnt we use rise over run and like put a point on 2 and go to the right 3 ?

OpenStudy (anonymous):

well \[ y = \frac{3}{2}x - 2 \] Since this is a linear function, the graph that comes out is a straight line. And all we need to draw this line are 2 points. You could take any 2 points you wish. let's say x=0 and x=1 \[ f(0) = \frac{3}{2} \cdot 0 -2 = -2 \\ f(1) = \frac{3}{2} \cdot 1 -2 = \frac{3}{2} - \frac{4}{2} = -\frac{1}{2}\\ \] And now just a ruler and to draw a line across them both.. In case you wanna find where it intersects with the X axis \[ y = 0 \\ \frac{3}{2}x - 2 = 0 \\ \frac{3}{2}x = 2 \\ 3x = 4 \\ x = \frac{4}{2} \] And for the intersection with Y axis.. we already have the result for x=0.

OpenStudy (anonymous):

is that correct ?

OpenStudy (anonymous):

Yes I would recommend this way: Since you know the y intercept plot that. Now make y = 0 to find the x intercept 2/3x-2=0 x=3 Plot (3,0): Now join the dots and you have your line

OpenStudy (anonymous):

thank you everyone